OpenStudy (anonymous):

To determine the number of deer in a game preserve, a conservationist catches 871 deer,tags them and lets them loose. Later 496 deer are caught; 248 of them are tagged. How many deer are in the preserve?

OpenStudy (bahrom7893):

1119... is that right?

OpenStudy (bahrom7893):

I meant if 496 deers are caught and 248 are tagged, then 496-248=248 are the ones that are not tagged..

OpenStudy (bobbyleary):

To determine the new total, you must find the "newly" tagged deer. 496 - 248 = 248 This now must be added to the original amount. 871 + 248 = 1119

OpenStudy (bahrom7893):

but maybe it's just: 871+496 = 1367

OpenStudy (anonymous):

I am not sure which one it would be?

OpenStudy (bahrom7893):

what i mean 248 of them are tagged.. he doesn't say tagged AGAIN, maybe he caught a bunch of untagged deers and tagged them all..

OpenStudy (bobbyleary):

Originally there were 871 caught. The on the second attempt, he caught a random group of 496 deer. 248 of these were already tagged from the past 871, but there are 496 - 248 = 248 new deer in the game preserve. So now add 871 + 248 to get the new total, or 1119.

OpenStudy (bahrom7893):

Bobby, the question doesn't say, were ALREADY tagged.. it says 496 are caught and 248 tagged.. probably our first thoughts were right, i got 1119 the same way, but ughh it's just the wording of the question is stupid..

OpenStudy (bahrom7893):

maybe they still didn't even catch all the deers haha..

OpenStudy (bahrom7893):

but yea prbbly 1119 is correct..

OpenStudy (anonymous):

yes 1119 is correct